OpenStudy (anonymous):

3b-13 +4b = 7b + 1 Please show all work-not just answer. I want to learn!!!

OpenStudy (zale101):

what does it mean when you combine like terms? Like, 3b, 4b, and 7b can be solved together but, can't be solved with -13 and 1. Because their not from the same species.

OpenStudy (anonymous):

So 14b -13 + 1 ?

OpenStudy (zale101):

i moved all the variables to one side to solve it 3b-7b+4b=13+1

OpenStudy (anonymous):

or did i dop

OpenStudy (anonymous):

that completely wrong....

OpenStudy (anonymous):

Yes I did that wrong haha. okay lemme see if i can solve that

OpenStudy (zale101):

you're a little close though but just follow along

OpenStudy (anonymous):

Okay yup I am going to need help solving this and thanks haha. :)

OpenStudy (anonymous):

8b = 13 + 1?

OpenStudy (zale101):